OpenStudy (twizttiez):

A container holds a combination of nickels and quarters. There are 156 coins in all, and their total value is $21.60. The system of equations models this situation, where x is the number of nickels and y is the number of quarters. x+y=156 0.05+0.25=21.6 How many nickels and quarters are in the container? A. 118 nickels and 38 quarters B. 106 nickels and 50 quarters C. 87 nickels and 69 quarters D. 72 nickels and 84 quarters

OpenStudy (twizttiez):

Can you help me @Fransisco_w ?

OpenStudy (anonymous):

Yea I'm just working the equation :D

OpenStudy (twizttiez):

Thank you

OpenStudy (anonymous):

I'm not exactly sure how to answer this, but this is what I always do. I usually start taking the answers and plugging them in... A. 0.05(118) + 0.25(38) = 21.6 5.9 + 9.5 = 21.6 I stop here because its obvious its no tthe answer B. 0.05 (106) + 0.25(50) = 21.6 5.3 + 12.5 = 21.6 Same as A. C. 0.5(87) + 0.25(69) = 21.6 4.35+17.25 = 21.6 21.6 = 21.6 Hurray its correct!

OpenStudy (anonymous):

Make sure to close the question and award medals. I hope this helped and have a nice day!

OpenStudy (twizttiez):

So C thanks i need help on a few more could you help i can open a new ? so you can get more medals

OpenStudy (anonymous):

Unfortunately I cannot D: I need to get to class. Create new posts for your other question and hopefully someone can help you.

OpenStudy (twizttiez):

Ok thanks anyways have a good day

OpenStudy (anonymous):

... its was C... I can help

OpenStudy (twizttiez):

One moment